How much of an alloy that is 20% copper should be mixed with 400 ounces of an alloy that is 70% copper in order to get an alloy that is 30% copper?

____ ounces

Answer by ankor@dixie-net.com(22740)

How much of an alloy that is 20% copper should be mixed with 400 ounces of an alloy that is 70% copper in order to get an alloy that is 30% copper?
:
Let x = amt of 20% copper is required:
Then
The resulting amt = (x+400)
:
.70(400) + .20(x) = .30(x+400)
:
280 + .2x = .3x + 120
:
.2x - .3x = 120 - 280
:
-.1x = -160
:
x = -160/-.1
:
x = +1600 oz of 20% copper is required
:
:
Check solution:
.7(400) + .2(1600) = .3(1600 + 400)
280 + 320 = .3(2000)
600 = 600; confirms out solution